OpenStudy (znappydooz):

Need some help in Algebra 2

OpenStudy (znappydooz):

What is the equation of the following graph in vertex form?

OpenStudy (znappydooz):

A: y = (x − 1)2 B: y = (x − 1)2 + 1 C: y = (x + 1)2 − 1 D: y = (x + 1)2

OpenStudy (mathmale):

The equation "in vertex form" is y-k=a(x-h)^2. Familiar with this?

OpenStudy (mathmale):

Please identify the coordinates of the vertex as showin in the illustration. Also please do the same with the y-intercept.

OpenStudy (znappydooz):

I dont get the vertex form @mathmale... Sorry, i had to do something else

OpenStudy (mathmale):

In the vertex form, you see y-k and x-h. If you're told that the vertex is at the point (-1,0), then h has the value -1 and k has the value 0. The letter a represents a constant multiplier which governs whether your parabola opens up widely or narrowly. (x,y) represents any point which is found on the graph.

OpenStudy (mathmale):

You are given the vertex, and also you're given one point that lies on the graph. Substitute these values into the vertex form I've given you (above). As before, (h,k) is the vertex; the illustration shows that the vertex is at which point? There's one other point shown on this graph. What are its coordinates? In other words, what are the values of x and y? If you can answer these questions, you'll have four numbers to substitute into the equation. The only quantity not given is "a," whose value you can find through algebra.

OpenStudy (mathmale):

If any of this is unclear to you, just ask questions for clarification.

OpenStudy (znappydooz):

(0)-(-1)=a(0-1)^2 Would that be the proper way to rewrite it?

OpenStudy (mathmale):

Hey, that looks good! Does this give you enuf info from which to find the value of 'a?'

OpenStudy (znappydooz):

Do I just solve that now?

OpenStudy (mathmale):

Yes. Try it. We'll need to check the value of 'a' we find to ensure that it's correct.

OpenStudy (znappydooz):

I ended up with a = 1

OpenStudy (mathmale):

So did I! Now please write the equation in vertex form. This time, do not substitute any values for x and y; just write x and y. (h,k) is (-1,0), as before.

OpenStudy (znappydooz):

Do you mean y-(-1)=a(x-1)^2 @mathmale?

OpenStudy (mathmale):

Remember that the vertex is at (-1,0). Therefore, h=-1 and k=0. So you should have y-0=(x-[-1])^2. Simplify that.

OpenStudy (mathmale):

So did I! Now please write the equation in vertex form. This time, do not substitute any values for x and y; just write x and y. (h,k) is (-1,0), as before.

OpenStudy (mathmale):

OpenStudy has been down, so my posts may appear out of order. Sorry. Sorry, too, that I need to get off the 'Net now. Good luck to you!